Adam who is 20 years old is 4 times as old as Mary what will be Mary's age when Adam is twice as old as her ?

Answer by fractalier(6550)

Mary must be 5 now. Let's call the number of years in the future, x.
Then we write
20 + x = 2(5 + x)
20 + x = 10 + 2x
x = 10 years
so that
Adam will be 30 and Mary will be 15.